# Gravitational Waves on Kerr Black Holes I: Reconstruction of Linearized Metric Perturbations

### v1

@@ D.2 Explicit metric components using only first derivatives
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(D^{mm}\mathop{}\!\partial_{r}+E^{mm}}\right)\overline{\hat{R}}^{(-2)}_{-\overline{\omega},\ell,-m}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)},$ (D.5b)
@@
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}}\Delta^{2}\mathopen{}\mathclose{{}\left(D^{mm}\mathop{}\!\partial_{r}+E^{mm}}\right)\hat{R}_{\omega\ell m}^{(+2)}\hat{S}_{\omega\ell m}^{(+2)},$ (D.7b)
@@ D.3 Explicit metric components with no derivatives at all
$H_{\omega\ell m}^{nm}$ $=-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\bigg{[}\mathopen{}\mathclose{{}\left(F^{nm}\overline{\hat{S}}_{-\omega,\ell,-m}^{(+2)}+G^{nm}\overline{\hat{S}}_{-\omega,\ell,-m}^{(-2)}}\right)\overline{\hat{R}}_{-\omega,\ell,-m}^{(+2)}$
$\phantom{=}\qquad\qquad+\mathopen{}\mathclose{{}\left(I^{nm}\overline{\hat{S}}_{-\omega,\ell,-m}^{(+2)}+J^{nm}\overline{\hat{S}}_{-\omega,\ell,-m}^{(-2)}}\right)\overline{\hat{R}}_{-\omega,\ell,-m}^{(-2)}\bigg{]},$ (D.9b)
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(F^{mm}\overline{\hat{R}}_{-\omega,\ell,-m}^{(+2)}+G^{mm}\overline{\hat{R}}_{-\omega,\ell,-m}}\right)^{(-2)}\overline{\hat{S}}_{-\omega,\ell,-m}^{(-2)},$ (D.9c)
@@
$F^{nn}$ $=\frac{\Delta^{3}\hat{\mathscr{C}}_{\omega\ell m}\hat{D}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}\overline{\mathbb{X}}_{-\overline{\omega},\ell,-m}},$ (D.10c)
$G^{nn}$ $=-\frac{\Delta^{3}\hat{\mathscr{C}}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{\overline{\mathbb{Y}}_{-\overline{\omega},\ell,-m}}{\overline{\mathbb{X}}_{-\overline{\omega},\ell,-m}}-\frac{a^{2}\sin{2\theta}}{\Sigma}}\right),$ (D.10d)
$I^{nn}$ $=\frac{\hat{D}_{\omega\ell m}}{\overline{\mathbb{X}}_{-\overline{\omega},\ell,-m}}\mathopen{}\mathclose{{}\left(\frac{\mathbb{D}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}+\frac{2r}{\Sigma}}\right),$ (D.10e)
$J^{nn}$ $=\frac{\overline{\mathbb{Y}}_{-\overline{\omega},\ell,-m}}{\overline{\mathbb{X}}_{-\overline{\omega},\ell,-m}}\mathopen{}\mathclose{{}\left(\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}+\frac{2r}{\Sigma}}\right)-\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.10f)
$F^{mm}$ $=-\frac{\Delta^{3}\mathscr{\hat{C}}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK-\Delta^{\prime}}{\Delta}+\frac{2}{\overline{\zeta}}}\right),$ (D.10g)
$G^{mm}$ $=\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK-\Delta^{\prime}}{\Delta}+\frac{2}{\overline{\zeta}}}\right)+\frac{6i\omega r+\lambda_{\omega\ell m}^{(-2)}}{\Delta}.$ (D.10h)
@@
$H_{\omega\ell m}^{ll}$ $=-\frac{\epsilon_{g}}{4}\zeta^{2}\mathopen{}\mathclose{{}\left(F^{ll}\hat{S}_{\omega\ell m}^{(+2)}+G^{ll}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(+2)},$ (D.11)
$H_{\omega\ell m}^{lm}$ $=\frac{\epsilon_{g}}{4\sqrt{2}}\frac{\zeta^{2}}{\overline{\zeta}}\Delta\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(F^{lm}\hat{S}_{\omega\ell m}^{(+2)}+G^{lm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(+2)}+\mathopen{}\mathclose{{}\left(I^{lm}\hat{S}_{\omega\ell m}^{(+2)}+J^{lm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(-2)}}\right],$ (D.12)
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}^{2}}\Delta\mathopen{}\mathclose{{}\left(F^{mm}\hat{R}_{\omega\ell m}^{(+2)}+G^{mm}\hat{R}_{\omega\ell m}^{(-2)}}\right)\hat{S}_{\omega\ell m}^{(+2)},$ (D.13)
where the new coefficients are
$F^{ll}$ $=-\frac{2\mathbb{W}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(Q+\cot{\theta}-\frac{ia\sin{\theta}}{\zeta}}\right)+6a\omega\cos{\theta}-\lambda_{\omega\ell m}^{(-2)},$ (D.14a)
$G^{ll}$ $=\frac{2\hat{D}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(Q+\cot{\theta}-\frac{ia\sin{\theta}}{\zeta}}\right),$ (D.14b)
$F^{lm}$ $=\mathopen{}\mathclose{{}\left(\frac{\mathbb{D}_{\omega\ell m}}{\mathbb{C}_{\omega\ell m}}-2\Delta^{\prime}}\right)\mathopen{}\mathclose{{}\left(\frac{\mathbb{W}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}-\frac{a^{2}\sin{2\theta}}{\Sigma}}\right)+\frac{\mathbb{W}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}\frac{2r\Delta}{\Sigma},$ (D.14c)
$G^{lm}$ $=\frac{\hat{D}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(iK+\frac{\mathbb{D}_{\omega\ell m}}{\mathbb{C}_{\omega\ell m}}}\right),$ (D.14d)
$I^{lm}$ $=\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}}{\mathbb{C}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{a^{2}\sin{2\theta}}{\Sigma}-\frac{\mathbb{W}_{\omega\ell m}}{\mathbb{Z}_{\omega\ell m}}}\right),$ (D.14e)
$J^{lm}$ $=\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}\hat{D}_{\omega\ell m}}{\mathbb{C}_{\omega\ell m}\mathbb{Z}_{\omega\ell m}},$ (D.14f)
$F^{mm}$ $=\mathopen{}\mathclose{{}\left(2\Delta^{\prime}-\frac{\mathbb{D}_{\omega\ell m}}{\mathbb{C}_{\omega\ell m}}}\right)\mathopen{}\mathclose{{}\left(\frac{2iK+\Delta^{\prime}}{\Delta}-\frac{2}{\zeta}}\right)-6i\omega r+\lambda_{\omega\ell m}^{(-2)},$ (D.14g)
$G^{mm}$ $=\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}}{\mathbb{C}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK+\Delta^{\prime}}{\Delta}-\frac{2}{\zeta}}\right).$ (D.14h)
@@ E.2 Functional form of the Wronskian
$w(z)\equiv z^{\gamma-1}(1-z)^{\delta-1}e^{\epsilon z},$ (E.6)
@@ F.1 Mapping the angular ODE to the confluent Heun equation
In this section, we closely follow Borissov and Fiziev [29] and map the angular ODE (1.11) the confluent Heun equation. First, we change to a variable $u=\cos{\theta}$, such that Eq. (1.11) becomes
@@ G.1 Mapping the radial ODE to the confluent Heun equation
In this section, we closely follow Borissov and Fiziev [29] and map the radial ODE (1.10) the confluent Heun equation. This equation has two regular singular points at $r=r_{\pm}$ and an irregular singular point at $r=\infty$. Once again, this means that via a change of variable and a field redefinition, it can be mapped to the canonical form (E.1) of the confluent Heun equation, for some parameters $\gamma$, $\delta$, $\epsilon$, $\alpha$, and $q$. To put it into this canonical form, we first transform
@@
$H^{\prime\prime}(r)+\mathopen{}\mathclose{{}\left(\frac{2\xi_{1}+s+1}{r-r_{+}}+\frac{2\xi_{1}+s+1}{r-r_{-}}+2\xi_{3}}\right)H^{\prime}(r)$
@@
$\xi_{2}=-ic_{-}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{+}}\right)\qquad\text{or}\qquad\xi_{2}=-s+ic_{-}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{+}}\right),$ (G.4b)
@@
In the absence of additional boundary conditions, Eq. (G.7) has two mode solutions, which take the simple forms (E.2) when expanded around $z_{\pm}=0$. If we choose the mapping $r\to z_{+}$, then these two modes will have a definite behavior at the outer horizon $z_{+}=0$. This choice results in the “in” and “out” radial modes (2.17).
@@ G.2 Derivation of the radial Teukolsky–Starobinsky constants
Ori [26] developed an elegant method for carrying out this computation. The idea is to exploit the known behavior (G.10) of the “in” and “out” modes near the outer horizon: [^24] Caveat lector: What we call the “in” and “out” modes, Ori calls “down” and “up” modes, and vice versa.
$\hat{R}_{\omega\ell m}^{(s)\,{\rm in}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}\Delta^{-s}e^{-ikr_{*}},$ (G.10a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm out}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}e^{ikr_{*}},$ (G.10b)
@@
It follows that at leading order near the outer horizon, we have in terms of $w=4Mkr_{+}$:
@@
$\mathscr{D}_{0}\mathopen{}\mathclose{{}\left(f(r)e^{-ikr_{*}}}\right)$ $\approx\mathopen{}\mathclose{{}\left(f^{\prime}(r)-\frac{iw}{\Delta}}\right)e^{ikr_{*}},$ (G.12b)
$\mathscr{D}_{0}^{\dagger}\mathopen{}\mathclose{{}\left(f(r)e^{ikr_{*}}}\right)$ $\approx\mathopen{}\mathclose{{}\left(f^{\prime}(r)+\frac{iw}{\Delta}}\right)e^{ikr_{*}},$ (G.12c)
@@
$\mathscr{D}_{0}^{4}\hat{R}_{\omega\ell m}^{(-2)\,{\rm in}}$ $\approx\mathscr{D}_{0}^{4}\mathopen{}\mathclose{{}\left(\Delta^{2}e^{-ikr_{*}}}\right)\approx\frac{\Gamma}{\Delta^{2}}e^{-ikr_{*}},$ (G.13a)
$\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\Delta^{2}\hat{R}_{\omega\ell m}^{(+2)\,{\rm out}}$ $\approx\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\mathopen{}\mathclose{{}\left(\Delta^{2}e^{ikr_{*}}}\right)\approx\widetilde{\Gamma}e^{ikr_{*}},$ (G.13b)
@@
Ori [26] showed that this method also works for the “up” and “down” modes, which we have ignored as we do not know how to represent them in terms of the function $\operatorname{HeunC}(z)$.14 However, such a representation is not needed here, as we only need their behavior at infinity,
$\hat{R}_{\omega\ell m}^{(s)\,{\rm up}}(r)$ $\stackrel{{\scriptstyle r\to\infty}}{{\approx}}\frac{e^{+i\omega r_{*}}}{r^{2s+1}},$ (G.14a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm down}}(r)$ $\stackrel{{\scriptstyle r\to\infty}}{{\approx}}\frac{e^{-i\omega r_{*}}}{r}.$ (G.14b)
@@
$\mathscr{D}_{0}^{4}\hat{R}_{\omega\ell m}^{(-2)\,\rm down}$ $\approx\mathscr{D}_{0}^{4}\mathopen{}\mathclose{{}\left(\frac{e^{-i\omega r_{*}}}{r}}\right)\approx 16\omega^{4}\frac{e^{-i\omega r_{*}}}{r},$ (G.17a)
$\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\Delta^{2}\hat{R}_{\omega\ell m}^{(+2)\,\rm up}$ $\approx\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\mathopen{}\mathclose{{}\left(\frac{e^{+i\omega r_{*}}}{r^{3}}}\right)\approx 16\omega^{4}\frac{e^{i\omega r_{*}}}{r}.$ (G.17b)

### v2

@@ D.2 Explicit metric components using only first derivatives
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(D^{mm}\mathop{}\!\partial_{r}+E^{mm}}\right)\hat{R}^{(-2)}_{\omega\ell m}\hat{S}_{\omega\ell m}^{(-2)},$ (D.5b)
@@
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}^{2}}\Delta^{2}\mathopen{}\mathclose{{}\left(D^{mm}\mathop{}\!\partial_{r}+E^{mm}}\right)\hat{R}_{\omega\ell m}^{(+2)}\hat{S}_{\omega\ell m}^{(+2)},$ (D.7b)
@@ D.3 Explicit metric components with no derivatives at all
$H_{\omega\ell m}^{nm}$ $=-\frac{\epsilon_{g}}{2\sqrt{2}\overline{\zeta}}\bigg{[}\mathopen{}\mathclose{{}\left(F^{nm}\hat{S}_{\omega\ell m}^{(+2)}+G^{nm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(+2)}+\mathopen{}\mathclose{{}\left(I^{nm}\hat{S}_{\omega\ell m}^{(+2)}+J^{nm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(-2)}\bigg{]},$ (D.9b)
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(F^{mm}\hat{R}_{\omega\ell m}^{(+2)}+G^{mm}\hat{R}_{\omega\ell m}^{(-2)}}\right)\hat{S}_{\omega\ell m}^{(-2)},$ (D.9c)
@@
$F^{nm}$ $=\frac{\Delta^{3}\hat{\mathscr{C}}_{\omega\ell m}\hat{D}^{\prime}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}\mathbb{X}_{\omega\ell m}},$ (D.10c)
$G^{nm}$ $=-\frac{\Delta^{3}\hat{\mathscr{C}}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{\mathbb{Y}_{\omega\ell m}}{\mathbb{X}_{\omega\ell m}}-\frac{a^{2}\sin{2\theta}}{\Sigma}}\right),$ (D.10d)
$I^{nm}$ $=-\frac{\hat{D}^{\prime}_{\omega\ell m}}{\mathbb{X}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{\mathbb{C}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}+\frac{2r}{\Sigma}}\right),$ (D.10e)
$J^{nm}$ $=\frac{\mathbb{Y}_{\omega\ell m}}{\mathbb{X}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}+\frac{2r}{\Sigma}}\right)-\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.10f)
$F^{mm}$ $=-\frac{\Delta^{3}\mathscr{\hat{C}}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK-\Delta^{\prime}}{\Delta}+\frac{2}{\zeta}}\right),$ (D.10g)
$G^{mm}$ $=\frac{\mathbb{B}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK-\Delta^{\prime}}{\Delta}+\frac{2}{\zeta}}\right)+\frac{6i\omega r+\lambda_{\omega\ell m}^{(-2)}}{\Delta}.$ (D.10h)
@@
$H_{\omega\ell m}^{ll}$ $=-\frac{\epsilon_{g}}{4}\zeta^{2}\mathopen{}\mathclose{{}\left(F^{ll}\hat{S}_{\omega\ell m}^{(+2)}+G^{ll}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(+2)},$ (D.11a)
$H_{\omega\ell m}^{lm}$ $=\frac{\epsilon_{g}}{4\sqrt{2}}\frac{\zeta^{2}}{\overline{\zeta}}\Delta\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(F^{lm}\hat{S}_{\omega\ell m}^{(+2)}+G^{lm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(+2)}+\mathopen{}\mathclose{{}\left(I^{lm}\hat{S}_{\omega\ell m}^{(+2)}+J^{lm}\hat{S}_{\omega\ell m}^{(-2)}}\right)\hat{R}_{\omega\ell m}^{(-2)}}\right],$ (D.11b)
$H_{\omega\ell m}^{mm}$ $=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}^{2}}\Delta\mathopen{}\mathclose{{}\left(F^{mm}\hat{R}_{\omega\ell m}^{(+2)}+G^{mm}\hat{R}_{\omega\ell m}^{(-2)}}\right)\hat{S}_{\omega\ell m}^{(+2)}.$ (D.11c)
Here, the new coefficients are
$F^{ll}$ $=\frac{2\mathbb{Y}_{-\omega,\ell,-m}}{\mathbb{X}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(Q+\cot{\theta}-\frac{ia\sin{\theta}}{\zeta}}\right)+6a\omega\cos{\theta}-\lambda_{\omega\ell m}^{(-2)},$ (D.12a)
$G^{ll}$ $=-\frac{2\hat{D}_{\omega\ell m}}{\mathbb{X}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(Q+\cot{\theta}-\frac{ia\sin{\theta}}{\zeta}}\right),$ (D.12b)
$F^{lm}$ $=\mathopen{}\mathclose{{}\left(\frac{\mathbb{C}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}+2\Delta^{\prime}}\right)\mathopen{}\mathclose{{}\left(\frac{\mathbb{Y}_{-\omega,\ell,-m}}{\mathbb{X}_{\omega\ell m}}+\frac{a^{2}\sin{2\theta}}{\Sigma}}\right)-\frac{\mathbb{Y}_{-\omega,\ell,-m}}{\mathbb{X}_{\omega\ell m}}\frac{2r\Delta}{\Sigma},$ (D.12c)
$G^{lm}$ $=-\frac{\hat{D}_{\omega\ell m}}{\mathbb{X}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(2\Delta^{\prime}-\frac{2r\Delta}{\Sigma}+\frac{\mathbb{C}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}}\right),$ (D.12d)
$I^{lm}$ $=-\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{a^{2}\sin{2\theta}}{\Sigma}+\frac{\mathbb{Y}_{-\omega,\ell,-m}}{\mathbb{X}_{\omega\ell m}}}\right),$ (D.12e)
$J^{lm}$ $=\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}\hat{D}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}\mathbb{X}_{\omega\ell m}},$ (D.12f)
$F^{mm}$ $=\mathopen{}\mathclose{{}\left(2\Delta^{\prime}+\frac{\mathbb{C}_{\omega\ell m}}{\mathbb{A}_{\omega\ell m}}}\right)\mathopen{}\mathclose{{}\left(\frac{2iK+\Delta^{\prime}}{\Delta}-\frac{2}{\zeta}}\right)-6i\omega r+\lambda_{\omega\ell m}^{(-2)},$ (D.12g)
$G^{mm}$ $=-\frac{\hat{\mathscr{C}}_{\omega\ell m}^{\prime}}{\mathbb{A}_{\omega\ell m}}\mathopen{}\mathclose{{}\left(\frac{2iK+\Delta^{\prime}}{\Delta}-\frac{2}{\zeta}}\right).$ (D.12h)
@@ E.2 Functional form of the Wronskian
$w(z)\equiv z^{\gamma-1}(z-1)^{\delta-1}e^{\epsilon z},$ (E.6)
@@ F.1 Mapping the angular ODE to the confluent Heun equation
In this section, we closely follow Borissov and Fiziev [30] and map the angular ODE (1.11) to the confluent Heun equation. First, we change to a variable $u=\cos{\theta}$, such that Eq. (1.11) becomes
@@ G.1 Mapping the radial ODE to the confluent Heun equation
In this section, we closely follow Borissov and Fiziev [30] and map the radial ODE (1.10) to the confluent Heun equation. This equation has two regular singular points at $r=r_{\pm}$ and an irregular singular point at $r=\infty$. Once again, this means that via a change of variable and a field redefinition, it can be mapped to the canonical form (E.1) of the confluent Heun equation, for some parameters $\gamma$, $\delta$, $\epsilon$, $\alpha$, and $q$. To put it into this canonical form, we first transform
@@
$H^{\prime\prime}(r)+\mathopen{}\mathclose{{}\left(\frac{2\xi_{1}+s+1}{r-r_{+}}+\frac{2\xi_{2}+s+1}{r-r_{-}}+2\xi_{3}}\right)H^{\prime}(r)$
@@
$\xi_{2}=-ic_{-}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{-}}\right)\qquad\text{or}\qquad\xi_{2}=-s+ic_{-}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{-}}\right),$ (G.4b)
@@
In the absence of additional boundary conditions, Eq. (G.7) has two mode solutions, which take the simple forms (E.2) when expanded around $z_{\pm}=0$. If we map $r\to z_{+}$, then we obtain the “in” and “out” radial modes (2.17), which have a definite behavior at the outer horizon $z_{+}=0$.
@@ G.2 Derivation of the radial Teukolsky–Starobinsky constants
Ori [26] developed an elegant method for carrying out this computation. The idea is to exploit the known behavior (2.20) of the “in” and “out” modes near the outer horizon: [^27] Caveat lector: What we call the “in” and “out” modes, Ori calls “down” and “up” modes, and vice versa.
$\hat{R}_{\omega\ell m}^{(s)\,{\rm in}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}c^{\rm in}\Delta^{-s}e^{-ikr_{*}},$ $c^{\rm in}=\mathopen{}\mathclose{{}\left[e^{ikr_{+}}\frac{\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{-ic_{-}k}}{\mathopen{}\mathclose{{}\left(r_{+}+r_{-}}\right)^{2ikM}}}\right],$ (G.10a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm out}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}c^{\rm out}e^{ikr_{*}},$ $c^{\rm out}=\mathopen{}\mathclose{{}\left[e^{-ikr_{+}}\frac{\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{ic_{-}k}}{\mathopen{}\mathclose{{}\left(r_{+}+r_{-}}\right)^{-2ikM}}}\right],$ (G.10b)
@@
It follows that at leading order near the outer horizon, we have in terms of $w=4Mkr_{+}$: [^28] This corrects Ori’s version of these equations, appearing between Eqs. (33) and (34), which are missing terms.
@@
$\mathscr{D}_{0}\mathopen{}\mathclose{{}\left(f(r)e^{-ikr_{*}}}\right)$ $\approx\mathopen{}\mathclose{{}\left(f^{\prime}(r)-\frac{iw}{\Delta}f(r)}\right)e^{-ikr_{*}},$ (G.12b)
$\mathscr{D}_{0}^{\dagger}\mathopen{}\mathclose{{}\left(f(r)e^{ikr_{*}}}\right)$ $\approx\mathopen{}\mathclose{{}\left(f^{\prime}(r)+\frac{iw}{\Delta}f(r)}\right)e^{ikr_{*}},$ (G.12c)
@@
$\mathscr{D}_{0}^{4}\hat{R}_{\omega\ell m}^{(-2)\,{\rm in}}$ $\approx c^{\rm in}\mathscr{D}_{0}^{4}\mathopen{}\mathclose{{}\left(\Delta^{2}e^{-ikr_{*}}}\right)\approx c^{\rm in}\frac{\Gamma}{\Delta^{2}}e^{-ikr_{*}},$ (G.13a)
$\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\Delta^{2}\hat{R}_{\omega\ell m}^{(+2)\,{\rm out}}$ $\approx c^{\rm out}\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\mathopen{}\mathclose{{}\left(\Delta^{2}e^{ikr_{*}}}\right)\approx c^{\rm out}\widetilde{\Gamma}e^{ikr_{*}},$ (G.13b)
@@
Ori [26] showed that this method also works for the “up” and “down” modes, which we have ignored as we do not know how to represent them in terms of the function $\operatorname{HeunC}(z)$.17 However, such a representation is not needed here, as we only need their behavior at infinity,
$\hat{R}_{\omega\ell m}^{(s)\,{\rm up}}(r)$ $\stackrel{{\scriptstyle r\to\infty}}{{\approx}}c^{\rm up}\frac{e^{+i\omega r_{*}}}{r^{2s+1}},$ (G.14a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm down}}(r)$ $\stackrel{{\scriptstyle r\to\infty}}{{\approx}}c^{\rm down}\frac{e^{-i\omega r_{*}}}{r}.$ (G.14b)
@@
$\mathscr{D}_{0}^{4}\hat{R}_{\omega\ell m}^{(-2)\,\rm down}$ $\approx c^{\rm down}\mathscr{D}_{0}^{4}\mathopen{}\mathclose{{}\left(\frac{e^{-i\omega r_{*}}}{r}}\right)\approx 16\omega^{4}c^{\rm down}\frac{e^{-i\omega r_{*}}}{r},$ (G.17a)
$\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\Delta^{2}\hat{R}_{\omega\ell m}^{(+2)\,\rm up}$ $\approx c^{\rm up}\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\mathopen{}\mathclose{{}\left(\frac{e^{+i\omega r_{*}}}{r^{3}}}\right)\approx 16\omega^{4}c^{\rm up}\frac{e^{i\omega r_{*}}}{r}.$ (G.17b)